if the areo of a given rectangle is 24 and the perimeter of the same rectangle is 22 . what are the measurements for the length and width?
Answer by Alan3354(69443)

if the areo of a given rectangle is 24 and the perimeter of the same rectangle is 22 . what are the measurements for the length and width?
--------------
P = 2L + 2W = 22
L + W = 11
L*W = 24
-------
Find 2 factors of 24 that add to 11.